INTERSTELLAR CHEMICAL MODELS WITH MOLECULAR ANIONS

\underline{ERIC HERBST}, Departments of Physics, Astronomy & Chemistry, The Ohio State University, Columbus, OH 43210; TOM MILLAR, MARTIN CORDINER, CATHERINE WALSH, Queen’s University Belfast, Belfast, Northern Ireland, UK; ROISIN NI CHUIMIN, Astrophysics Group, The University of Manchester, Manchester M60 1QD, UK.

Now that the first molecular anion, C_{6}H^{−}, has been detected in the cold dark cloud TMC-1 and both this ion and C_{4}H^{−} detected in the circumstellar envelope IRC+10216 based on the laboratory work of Thaddeus and co-workers^{a}^{b}, it is of interest to study anew how anions are produced and destroyed via gas-phase reactions in interstellar and circumstellar environments. Relatively abundant anions are most likely formed principally by radiative attachment and depleted by photodetachment and ion-molecule reactions with atomic hydrogen and oxygen^{c}. New estimates of rate coefficients for some of these processes have been made. We have then included formation and depletion pathways for a number of molecular anions in the latest UMIST database for astronomy (http://www.udfa.net/). With this enhanced network, theoretical abundances and column densities of the anions in TMC-1 and IRC+10216 have been calculated. Our results will be discussed, along with our views on the most likely new anions to be detected once laboratory data become available.

aM. C. McCarthy et al. ApJ \underline{652}, L141, 2006.

bH. Gupta et al. ApJ \underline{655}, L57, 2007.

cE. Herbst, Nature \underline{289}, 656, 1981.